# Mahalanobis k-NN: A Statistical Lens for Robust Point-Cloud Registrations

### v1

@@ 5.1 Point Cloud Registration
We build upon the model architecture, training strategies, evaluation metrics, and hyperparameters established in DCP [DCP]  [^1] https://github.com/WangYueFt/dcp. We introduce the Mahalanobis version of DCP (MDCP), detailed in Section 4.1. MDCP comes in two variants: v1 without a transformer and v2 with a transformer. We further extend our approach to DeepUME [DeepUME] by proposing the Mahalanobis variant, MDeepUME. For a fair comparison, we benchmark geodesic versions of these models using vectorized Floyd-Warshall on Euclidean k-NN graphs.
Dataset. To facilitate benchmarking, we employ publicly available datasets: ModelNet40 [modelnet], FAUST [FAUST], and Stanford3D (S-3D) [^2] http:// graphics.stanford.edu/data/3Dscanrep/ . Notably, the latter two datasets are exclusively employed for testing purposes. To ensure methodological consistency, we adhere to the official training/testing splits and settings as stipulated in the original works of DCP [DCP] and DeepUME [DeepUME]. Among these datasets, our approach is rigorously evaluated.
@@ 5.2 Comparison with State-of-the-art Methods
Table 1: Comparison on ModelNet40 [modelnet] for unseen data, unseen category, and Gaussian noise. MDCP-(v1,v2) and MDeepUME outperform baselines (bold best, underline second best). Lower is better. $\dagger$ denotes reproduced results, -Geo refers to Floyd-Warshall geodesic-graph version.
@@
ModelNet40 Results. Table 1 compare our methods (MDCP, MDeepUME) with baselines and their geodesic versions. We outperform all methods across various settings: Unseen data: Train-test split from ModelNet40 [modelnet]; Unseen category: First 20 classes for training, remaining for testing; Gaussian Noise: $\mathcal{N}(0,0.01)$ clipped to [-0.05, 0.05]. The point density is 1024 in all settings. Baseline evaluations are from [DCP] (highly comparable results achieved with minimal standard deviation during our reproduction attempts). For each case, a rigid transformation is applied along each axis (rotation: $[0^{\circ},45^{\circ}]$, translation: [-0.5, 0.5]). The network receives both the original and transformed point cloud as input, evaluated against ground truth using MDCP and MDeepUME.
Figure 3: Robustness evaluation on unseen data (FAUST [FAUST] hand), low-density source (airplane), and low-density target (chair) point clouds (trained on ModelNet40 [modelnet]). Red boxes highlights vulnerabilities in DCP-v1 [DCP], while green boxes highlights MDCP-v1’s efficacy.
Table 2: Evaluation of few-shot classification on ModelNet40 [modelnet] and ScanObjectNN [SONN] (OBJ ONLY, OBJ+BG, PB75 splits). Our DCGNN in MDCP-(v1,v2) outperforms baselines across all settings (bold denotes best accuracy). We report mean accuracy (%) and standard deviation from 50 runs.
ModelNet40 [modelnet]
@@
Few-shot Classification. Table 2 evaluates the discriminative power of Mahalanobis distance. We compare features learned by DCP-v1/v2 [DCP] and MDCP-v1/v2 in few-shot classification tasks using the $k$-way, $m$-shot setting [CrossPoint]. Models are trained for registration on ModelNet40 [modelnet] and then tested on few-shot classification for ModelNet40 and three ScanObjectNN [SONN] splits (OBJ ONLY, OBJ+BG, PB75).
MDCP consistently outperforms DCP across all settings. Summarizing Table 2, MDCP-(v1,v2) achieves average accuracy improvement of $(22.98,6.52)$ on ModelNet40. For ScanObjectNN, improvements are: OBJ ONLY: $(12.68,8.19)$; OBJ+BG: $(25.99,18.40)$; PB75: $(25.06,17.11)$, [(a, b) denotes average accuracy increment (%) for MDCP-(v1,v2)].
These results suggest two key points: 1) v1 (without transformer) outperforms v2, indicating transformers may struggle with limited data; 2) Even on unseen ScanObjectNN data, MDCP shows consistent improvement across splits.
@@ 5.3 Endurance Test
Table 3: Quantitative comparison of MDCP-(v1,v2) robustness against original on ModelNet40 [modelnet] for low-density point cloud registration. We evaluate across various settings. MDCP consistently outperforms its original counterpart (bold denotes best). Lower is better for RMSE(t) (reported in $10^{-2}$ units). Note: -Geo refers to the vectorized Floyd-Warshall geodesic-graph version.
Target Low Density ModelNet40 [modelnet]
@@
Source Low Density ModelNet40 [modelnet]
1024 DCP-v1 [DCP] 43.2637 1.1146 8.304669 1.228 38.6104 1.0306
@@
Efficiency towards Point-density. Table 3 evaluates MDCP’s robustness against varying point densities, simulating real-world sensor data [survey]. One point cloud (source or target) is downsampled by half (e.g., 2048 → 1024). Evaluation metrics use the same number of points for both clouds (e.g., source with 1024 points and target with 2048 points are used to compute transformation, then applied to original points for error calculation).
Figure 3 compares DCP-v1 [DCP] and MDCP-v1. When densities are similar (man’s face), both perform well. However, MDCP-v1 demonstrates superior robustness under varying densities (chair). This highlights its effectiveness in real-world applications where source or target point clouds may have different densities due to factors like sensor capabilities or data acquisition methods (e.g., SFM-LiDAR fusion).
@@
Table 4: Quantitative comparison of proposed MDeepUME’s robustness on ModelNet40 [modelnet], FAUST [FAUST], and Stanford 3D datasets. We evaluate across various noise, benchmarked by authors in DeepUME [DeepUME]. MDeepUME consistently outperforms DeepUME (bold denotes best results). Lower is better for all metrics: RMSE(t), chamfer distance (CD) [ChamferDistance], and Hausdorff distance (HD) [HausdorffDistance] (reported in $10^{-2}$ units).
Noise-types DeepUME [DeepUME] MDeepUME
@@
Efficiency towards Various Noise Types. Table 4 evaluates the robustness of MDeepUME against various noise types from DeepUME [DeepUME] (Bernoulli, Gaussian, Sampling, Zero-Intersection). We benchmark on ModelNet40 [modelnet] (training and evaluation), FAUST [FAUST], and Stanford3D (testing only). Training settings follow the original paper [DeepUME] ( [^3] https://github.com/langnatalie/DeepUME).
MDeepUME outperforms DeepUME across datasets and noise types in terms of Chamfer distance (CD) [ChamferDistance], Hausdorff distance (HD) [HausdorffDistance], and root mean squared error in rotation (RMSE(R)). However, limitations are observed in translation accuracy (RMSE(t)) due to DeepUME’s inherent projection into SO(3)-invariant space, where Mahalanobis distance is less effective.
@@ 6 Conclusion
Our paper introduces Mahalanobis k-NN, a statistical framework for improving point cloud registration in learning-based methods. It addresses challenges like feature matching due to variations in point cloud densities. Mahalanobis k-NN leverages local neighborhood distribution for accurate feature extraction, outperforming methods like Flyod-Warshall. It integrates seamlessly into local graph-based point cloud analysis methods like DCP and DeepUME, achieving state-of-the-art performance on benchmark datasets. We also demonstrate that registered point cloud features possess discriminative capabilities, leading to significant improvements in few-shot classification tasks. Mahalanobis k-NN offers exceptional surface awareness and outlier robustness, benefiting tasks like segmentation, decomposition, and normal estimation. Additionally, it serves as a surficial geometry-based error metric, enhancing tasks like generation and upsampling. Its versatility makes it a crucial component in advancing point cloud processing, warranting further benchmarking to highlight its significance.
[…]

### v2

@@ 5.1 Point Cloud Registration
We build upon the model architecture, training strategies, evaluation metrics, and hyperparameters established in DCP [42]  [^1] https://github.com/WangYueFt/dcp. We introduce the Mahalanobis version of DCP (MDCP), detailed in Section 4.1. MDCP comes in two variants: v1 without a transformer and v2 with a transformer. We further extend our approach to DeepUME [27] by proposing the Mahalanobis variant, MDeepUME. For a fair comparison, we benchmark geodesic versions of these models using vectorized Floyd-Warshall on Euclidean k-NN graphs.
Dataset. To facilitate benchmarking, we employ publicly available datasets: ModelNet40 [45], FAUST [7], and Stanford3D (S-3D) [^2] http:// graphics.stanford.edu/data/3Dscanrep/ . Notably, the latter two datasets are exclusively employed for testing purposes. To ensure methodological consistency, we adhere to the official training/testing splits and settings as stipulated in the original works of DCP [42] and DeepUME [27]. Among these datasets, our approach is rigorously evaluated.
@@ 5.2 Comparison with State-of-the-art Methods
Table 1: Quantitative comparison of state-of-the-art methods benchmarked on ModelNet40 [45] Dataset for three tasks 1) Unseen data, 2) Unseen category, 3) Robustness towards Gaussian Noise; Our proposed MDCP-(v1,v2) and MDeepUME; outperforms its original variants in almost all cases. All evaluation metrics are lower the better. We highlight bold as best and underline as second best. Note: $\dagger$ depicts results reproduced on our systems and -Geo refers to Flyod-Warshall geodesic-graph version.
@@
Results on ModelNet40 Dataset [45]. Comprehensive results are presented in Table 1, offering a thorough comparison between our method, state-of-the-art alternatives and their geodesic versions. The presented outcomes highlight the superior performance of our method across all evaluation metrics and in diverse settings. Our benchmarking encompasses scenarios including Unseen-data evaluation, where the train-test split outlined in ModelNet40 [45] is utilized; Unseen category evaluation, where the first 20 classes of ModelNet40 are employed for training and the remaining for testing; and a robustness assessment against Gaussian Noise $\mathcal{N}(0,0.01)$, which is clipped to the interval [-$0.05,0.05$] before being added to the point cloud. Notably, a point density of 1024 is consistently sampled in all aforementioned settings. Note: Our assessment of state-of-the-art methods is derived from [42], as our efforts to reproduce the open-source code yielded highly comparable results with minimal standard deviation.
In each case, a rigid transformation is introduced along each axis, with rotation uniformly sampled in the range [$0^{\circ},45^{\circ}$] and translation within [-$0.5,0.5$]. The input to the network consists of both the original point cloud and the transformed point cloud through the rigid motion. This input is subsequently evaluated against the known ground truth in both our proposed MDCP and MDeepUME methods.
Figure 3: Evaluation of Robustness for varied test cases. Both DCP-v1 [42] and proposed MDCP-v1 trained on ModelNet40 [45] dataset. Our evaluation encompasses three distinct scenarios: 1) Evaluation on unseen data, specifically a man holding palms together from the FAUST [7] dataset; 2) Assessment of low-density source point clouds, exemplified by an airplane; and 3) Examination of low-density target point clouds, illustrated by a chair. In our visual analysis, regions highlighted in red highlights the vulnerabilities observed in vanilla DCP-v1 [42], while regions in green accentuate the pronounced efficacy demonstrated by the proposed MDCP-v1 over the conventional DCP-v1.
Figure 4: We present a compelling demonstrations showcasing the surface awareness within the proposed MDCP-v1, in contrast to the Euclidean approach in DCP-v1[42]. Each row within the illustration corresponds to K values ranging from 2 to 5, effectively exemplifying the pronounced surficial awareness embodied by MDCP-v1. This surficial awareness is pivotal for robust point cloud registration tasks.
Furthermore, to substantiate the assertion that Mahalanobis distance operates as a statistical lens for enhanced feature matching, we undertake KMeans [20] clustering within the initial embedding space of DGCNN. This embedding space, a 64-dimensional domain following the first edge-convolution [44], is employed both in DCP-v1 and MDCP-v1. Our findings are presented visually through Figure 4. The depicted results unequivocally showcase the preservation of surface awareness in the case of the Mahalanobis variant, even in scenarios involving overclustering (e.g., K=5). This outcome stands in contrast to its Euclidean counterpart. Particularly is the consistent maintenance of surface awareness throughout various clustering cases, particularly evident in the context of the Airplane dataset. This affirmation highlights that the Mahalanobis version of DCP learns to discern and emphasize surface-aware regions that are crucial in achieving precise feature matching.
Table 2: We present the outcomes of few-shot classification experiments conducted on the ModelNet40 [45] dataset, as well as on three distinct splits (OBJ ONLY, OBJ+BG, and PB75) of the ScanObjectNN [41] dataset. Our proposed DCGNN employed in MDCP-(v1,v2) consistently surpasses its original variants by a substantial margin across all few-shot settings and datasets. The reported quantitative results in terms of accuracies (%) are presented below, with values in bold indicating the best performance. Additionally, we provide the standard deviation values derived from 50 runs for a comprehensive understanding of the results.
ModelNet40 [45]
@@
Few-shot Evaluation. The discriminative prowess of the proposed Mahalanobis distance is evaluated through Table 2, illustrating the performance of features learned by both DCP-(v1,v2) [42] and our proposed MDCP-(v1,v2) in the context of few-shot evaluation tasks. The evaluation adheres to a $k$-way, $m$-shot few-shot strategy detailed in  [1]. The models are initially trained for the point-cloud registration task on the ModelNet40 dataset, and subsequently assessed for few-shot classification tasks on the ModelNet40 test set. Furthermore, evaluations extend to three distinct splits, namely OBJ ONLY, OBJ + BG, and PB75, from the ScanObjectNN [41] dataset.
A striking observation is the substantial performance enhancement of the Mahalanobis version of DCP, which surpasses its original variant by a significant margin across all few-shot tasks. Summarizing the performance of Mahalanobis versions (v1,v2) as reported in Table 2, we increment across all few-shot tasks on the ModelNet40 [45] dataset of (22.98, 6.52). In the context of the ONLY OBJ split, an average increment of (12.68, 8.19) is recorded, while for the OBJ + BG split, the increments are (25.99, 18.40). Similarly, for the PB75 split, increments of (25.06, 17.11) are documented. Note: (a, b) denotes average accuracy increment in % for MDCP-(v1,v2).
From the results, two insights emerge: Firstly, v1 without a transformer outperforms v2 with a transformer, suggesting that transformers may not generalize as effectively with limited data. Secondly, its worth noting that even when evaluated on unseen ScanObjectNN [41] data, which was not included during the training phase, a consistent increment is observed across all splits.
@@ 5.3 Endurance Test
Table 3: Quantitative comparison to evaluate the robustness of the proposed MDCP-(v1,v2) against its original variants, benchmarked on the ModelNet40 [45] dataset. This comparison is conducted across various settings for the point-cloud registration task, particularly when the target and source point cloud exhibit low point density. Throughout the comparison, our approach consistently demonstrates evident superiority in performance. The bold typeface denotes the best results. It is important to note that all evaluation metrics follow a “lower is better” criterion, and the RMSE(t) metric is reported in units of $10^{-2}$. Note: -Geo refers to Flyod-Warshall geodesic-graph version.
Target Low Density ModelNet40 [45]
@@
Source Low Density ModelNet40 [45]
1024 DCP-v1 [42] 43.2637 1.1146 8.304669 1.228 38.6104 1.0306
@@
Efficiency towards Point-density. Table 3 evaluates MDCP’s robustness against varying point densities, simulating real-world sensor data [23]. One point cloud (source or target) is downsampled by half (e.g., 2048 → 1024). Evaluation metrics use the same number of points for both clouds (e.g., source with 1024 points and target with 2048 points are used to compute transformation, then applied to original points for error calculation).
Figure 3 compares DCP-v1 [42] and MDCP-v1. When densities are similar (man’s face), both perform well. However, MDCP-v1 demonstrates superior robustness under varying densities (chair). This highlights its effectiveness in real-world applications where source or target point clouds may have different densities due to factors like sensor capabilities or data acquisition methods (e.g., SFM-LiDAR fusion).
@@
Efficiency towards Point-density. Table 3 illustrates the robustness of the proposed Mahalanobis version of DCP [42] in the context of point-cloud registration tasks that simulate real-world scenarios. These scenarios involve two point clouds originating from distinct sensors [23], potentially exhibiting differing point densities. In this experimental setup, one of the two point clouds, either the source or target, is downsampled by half relative to the other. As depicted in Table 3, the value 1024 signifies that either the target or source point cloud is downsampled to half its original size (e.g., 2048 $\to$ 1024). Note: To compute evaluation metrics, both source and target point clouds are ensured to have the same number of points. Thus, rigid-transformation parameters are computed using the downsampled point cloud and applied to the original-scale points (e.g., $X_{2048}$ and $Y_{1024}$ are used to generate transformation parameters $g$, which are then utilized to evaluate RMSE($X_{2048}$, $g(Y_{2048})$)).
This scenario is effectively illustrated in Figure 3, where the performance comparison between DCP-v1 [42] and MDCP-v1 is particularly insightful. In cases where the point clouds have same density, such as in the highlighted regions depicting a man’s face, DCP-v1 and MDCP-v1 exhibit relatively close performance. However, the subtleties become more pronounced when subjected to the aforementioned settings. Specifically, DCP-v1 struggles to accurately estimate rigid transformations, leading to subtle angular discrepancies as evident from the highlighted regions. In contrast, the proposed MDCP-v1 showcases a robustness to varying point densities, as demonstrated through the highlighted regions portraying a chair.
Note: The source point cloud is a transformed version of $g(Y)$, wherein the objective is to estimate the rigid transformation parameters $g$ that align it perfectly with the target point cloud $X$.
Table 4: Quantitative comparison to evaluate the robustness of the proposed MDeepUME against its original variant, benchmarked on the ModelNet40 [45], FAUST [7] and Stanford 3D repository dataset. This comparison is conducted across various noise types proposed in DeepUME [27] for the point-cloud registration task. Throughout the comparison, our approach consistently demonstrates evident superiority in performance. The bold typeface is utilized to denote the best results. It is important to note that all evaluation metrics follow a “lower is better” criterion, and the RMSE(t), chamfer-distance [3] (CD), and hausdorff-distance [14] (HD) metric are reported in units of $10^{-2}$.
Noise-types DeepUME [27] MDeepUME
@@
Efficiency towards Various Noise Types. Table 4 illustrates the resilience of the proposed Mahalanobis version of DeepUME [27] in the context of point cloud registration tasks subjected to various noise sampling strategies, as outlined in DeepUME [27]. Our evaluation encompasses diverse noise types, including Bernoulli, Gaussian, Sampling, and Zero-Intersection, each meticulously detailed within the framework of DeepUME [27]. The benchmarking is conducted on the ModelNet40 [45], FAUST [7], and Stanford-3D Repository datasets. While the latter two datasets are exclusively utilized for testing, ModelNet40 is employed for comprehensive evaluation. All training-testing settings [^3] https://github.com/langnatalie/DeepUME are considered wrt original paper [27].
The outcomes presented in Table 4 substantiate the superior performance of the proposed MDeepUME over its original counterpart across various datasets and noise types. This observation holds true for a multitude of evaluation metrics, notably including Chamfer distance [3] (CD), Hausdorff distance [14] (HD), and root mean squared error in rotation (RMSE(R)). However, it is worth noting that MDeepUME demonstrates certain limitations, particularly in estimating translation accuracy, as evident from the RMSE(t) values in Table 4. This outcome is attributed to DeepUME’s intrinsic projection of points into an $\mathcal{SO}(3)$-invariant space, wherein the potential of Mahalanobis distance cannot be fully harnessed.
5.4 Limitations
While our approach harnesses the prowess of Mahalanobis distance for enhanced feature matching and point cloud registration, it is important to acknowledge its limitations. Mahalanobis distance computation critically depends on the accurate estimation of the covariance matrix. When the covariance matrix is ill-conditioned, the precision may become compromised. Although we mitigate this concern by introducing a small bias term ($10^{-5}$) to the diagonal of the covariance matrix. This limitation is evident in noise benchmark evaluation for DeepUME, where the Mahalanobis version fails in accurately estimating translation parameters. Despite these limitations, our approach significantly contributes to address key challenges in point cloud registration.
@@ 6 Conclusion
Our paper introduced Mahalanobis k-NN, a statistical framework for improving point cloud registration in learning-based methods. It addresses challenges like feature matching due to variations in point cloud densities. Mahalanobis k-NN leverages local neighborhood distribution for accurate feature extraction, outperforming methods like Flyod-Warshall. It integrates seamlessly into local graph-based point cloud analysis methods like DCP and DeepUME, achieving state-of-the-art performance on benchmark datasets. We also demonstrated that registered point cloud features possess discriminative capabilities, leading to significant improvements in few-shot classification tasks. Mahalanobis k-NN offers surface awareness and outlier robustness. Overall, our proposed method offers a comprehensive and effective solution for point cloud registration, providing superior results, and versatility across various benchmarking test-beds.